Suppose you wanted to use Earth’s magnetic field to make an ac generator at a location where the magnitude of the field is 0.05000 mT. Your coil has 1008 turns and a radius of 3.700 cm. At what angular velocity would you have to rotate it in order to generate an emf of amplitude 1.000 V?
Answer #1

given

magnetic field B = 0.05*10^-3 T

N = 1008 turns

radius r = 0.037 m

voltage V = 1 V

from the relation

angular velocity w = V/NBA

w = 1/1008*0.05*10^-3*3.14*0.037^2

w = 4615.7 rad/s